The area of the rectangle is 121 square feet.if it's width is 7 and one third feet,find its length
Answer by amarjeeth123(569)

Let the length be x.
Width=7 and 1/3 ft=22/3 ft
Area=121 square feet
Area=length*width
121=22/3*x
x=(3*121)/22=16.5 feet
The length is 16.5 feet.